40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12952     Entry time: Thu Apr 27 16:41:13 2017
Author: Eric Gustafson 
Type: Update 
Category: LSC 
Subject: Status of the 40 m PD Frequency Response Fiber System 
There two reports in the DCC describing the state of the system as of October 2014 including: (1)  Alex Cole’s “T1300618 Automated photodiode Frequency Response Measurement System” and a Wiki  created by Alex Cole where there are some instructions on the Master Script at https://wiki.ligo.caltech.edu/ajw?AlexanderCole    

And (2)  P140021 “Final Report: Automated Photodiode Frequency Response Measurement System for 40m Lab” by Nichin Sreekantaswamy and also as part of Nichin’s report by there is an archive of data at   https://wiki-40m.ligo.caltech.edu/Electronics/PDFR%20system   

I made a visual inspection of the system and saw that the following fibers collimators are still mounted in alignment mounts and the fiber is attached and pointed at a photodetector but possibly not aligned. 

ASP Table

Photodetector Label                             Fiber Label

REFL11                                              REFL55 Fiber on mount        

REFL33                                              REFL33 Fiber on mount

REFL55                                              REFL11 Fiber on mount

REFL165                                            No Fiber

AS55                                                   AS55 Fiber on mount

MCREFPD                                         MCREFPD Fiber on mount

No PD                                                 Loose unlabeled Fiber No mount

 

ITMX Optics Table

Photodetector Label                             Fiber Label

POX11                                                POX11 on mount

Unlabeled PD                                      POP22/POP110 on mount

NO PD                                                 POP55 loose fiber No mount

 

The RF switch seems to be hooked up and there is a fiber running from the Diode Laser module to the fiber splitter module. So REFL 11 and REFL545 seem to be illuminated by the wrong fiber. I’ll try and run the software on Monday and check to see if I need to move the fibers or just relabel them.

ELOG V3.1.3-